Operation Summary for May 5, 2010 - May 7, 2010
Wednesday May 5
The day shift began with Operations monitoring the Tevatron (TeV) store 7793 (256.13E30) and the Recycler stash (72.1E10), with the Antiproton Source (Pbar) stacking (12.7E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI, MiniBooNE, and Meson MTest experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.
During the day shift, Operations disabled the MTest beam for an access. Operations disabled the NuMI beam at 12:25 PM for an access to investigate the helium leak at the target. Stacking was down for an hour while Pbar personnel worked on a Debuncher kicker power supply (D:IKIK). Operations restored beam to MTest at 1:25 PM. Operations restored beam to NuMI at 3:10 PM.
During the evening shift, Operations disabled the MTest beam at 6:25 PM.
Thursday May 6
The midnight shift began with Operations monitoring store 7793 (67.03E30) and the stash (388.0E10), with Pbar stacking (10.0E10), and with NuMI and MiniBooNE taking beam.
During the midnight shift, Operations terminated store 7793 at 1:35 AM, put the TeV into HEP shot setup and established store 7794 at 2:40 AM with an average initial luminosity (AIL) of 364.0E30. Operations established beam to MTest at 4:25 AM.
During the day shift, MTA personnel conducted beam studies. NuMI personnel performed target scans. The Pbar Accumulator extraction Lambertson (D:ELAM) tripped off and required a power supply control card replacement. Preaccelerator personnel accessed the I- Source for repairs. A MI Lambertson (LAM52) began causing problems at the end of the shift.
During the evening shift, NuMI personnel completed their target scans at 4:30 PM. EE Support personnel adjusted the regulation for LAM52. Controls personnel replaced a 5-volt power supply for a MI readback system. TeV personnel conducted an end-of-store crystal collimator study from 8:30 to 11:30 PM. Operations terminated store 7794 at 11:45 PM and put the TeV into shot setup.
Friday May 7
The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (432.4E10), with Pbar stacking (41.6E10), and with NuMI and MiniBooNE taking beam.
During the midnight shift, Operations established store 7795 at 1:25 AM with an AIL of 397.0E30. Store 7795 quenched 2 hours later when a TeV antiproton abort kicker (#5) prefired causing quenches at TeV sectors A1, A2, A4, B1, F4, and to the B0 and D0 low beta magnets. By the end of shift, operators had prepared D0, TeV A-E, Enclosure B, and Booster for accesses.
The Plans for Friday and the Weekend
The Run Coordinator’s plan for today is to allow D0 to conduct a major repair to their silicon detector, which will take about 10 hours. The Run Coordinator will also allow the following work: TeV personnel will pull out their crystal collimator, Booster personnel will repair RF stations and replace a power amplifier, cryo system personnel will replace the D1 cold compressors, CDF will conduct dry engine work and other miscellaneous jobs. The plans for the weekend are to stack, stash, and store.
Comments and Suggestions
|last modified 05/07/2010 email Fermilab|
|Security, Privacy, Legal|